OpenStudy (anonymous):

How to convert 3pi/4 rad to degree, minute, and seconds (° ' ")

OpenStudy (shamim):

\[\pi rad= 180 degree\]

OpenStudy (shamim):

so\[\frac{ 3 \pi }{ 4 }=\frac{ 3 \times 180 }{ 4 }\]